# A randomized small-block Lanczos method for large-scale null space computations

### v1

@@ Abstract
Computing the null space of a large sparse matrix $A$ is a challenging computational problem, especially if the nullity—the dimension of the null space—is large. When using a block Lanczos method for this purpose, conventional wisdom suggests to use a block size $d$ that is not smaller than the nullity. In this work, we show how randomness can be utilized to allow for smaller $d$ without sacrificing convergence or reliability. Even $d=1$, corresponding to the standard single-vector Lanczos method, becomes a safe choice. This is achieved by using a small random diagonal perturbation, which moves the zero eigenvalues of $A^{\mathsf{T}}A$ away from each other, and a random initial guess. We analyze the effect of the perturbation on the attainable quality of the null space and derive convergence results that establish robust convergence for $d=1$. As demonstrated by our numerical experiments, a smaller block size combined with restarting and partial reorthogonalization results in reduced memory requirements and computational effort. It also allows for the incremental computation of the null space, without requiring a priori knowledge of the nullity.
@@ 1 Introduction
When $A$ is large and has no particular structure beyond being sparse, one needs to resort to Krylov subspace methods, which utilize matrix-vector products with $A$ and its transpose $A^{\mathsf{T}}$. Let $N$ denote the null space dimension of $A$. When $N=1$, standard choices include the Lanczos method for computing the smallest eigenvalue and eigenvector of $A^{\mathsf{T}}A$ or the mathematically equivalent, but numerically more reliable Golub–Kahan bidiagonalization directly applied to $A$ [10, Chapter 10]. When $N>1$, the Lanczos method (with a single starting vector) becomes unreliable and prone to miss components of the null space. In fact, in exact arithmetic only a one-dimensional subspace of the null space can be extracted via Lanczos [23, Section 12.2]. Block Lanczos methods address this issue by utilizing a block of $d>1$ starting vectors instead of a single starting vector [23, Section 13.10]. Common wisdom (see [23, Section 13.12] for examples) and existing convergence analyses all suggest to choose $d$ not smaller than the size of the eigenvalue cluster of interest. Applied to null space computation this would require $d\geq N$ and in exact arithmetic this condition is indeed necessary. The presence of round-off error blurs the situation. As an example, consider the $420\times 420$ diagonal matrix
(1) $A=\mathrm{diag}(0,\dotsc,0,1,2,\dotsc,399),$
that is, the first $21$ diagonal entries are zero. When applying block Lanczos to $A$ in order to compute its null space and choosing a block of $d$ (Gaussian) random starting vectors, one would expect to obtain no more than $d$ (approximate) zero eigenvalues and corresponding eigenvectors. However, when executing block Lanczos in double precision arithmetic without breakdown and declaring eigenvalues less than $10^{-4}$ as zero, we obtained the results reported in Table 1.
Table 1: Null space dimension obtained when applying block Lanczos with $d$ random starting vectors to the matrix from Eq. 1.
Block size $d$ 1 2 3 4 5 6 7 10 12 14 15 20 21
Dimension 11 11 16 12 15 15 21 20 21 21 21 21 21
While the null space dimension estimated by block Lanczos for small values of $d$ is erratic and smaller than $21$, it is always larger than $d$. In fact, already for $d=12$, the correct null space dimension is detected. Experiments with other matrices lead to similar findings, suggesting that round-off error breaks some of the curse incurred by eigenvalue clusters, but it does not fully address the convergence issues of single-vector or small-block Lanczos methods either.
@@
Inspired by [19], we propose a randomized small-block Lanczos method for null space calculation, outlined in Algorithm 1. Compared to the task of low-rank approximation, there are two major differences. First, in low-rank approximation, the largest few singular values are usually unknown and, in most applications, they do not form tight large clusters. In the context of null spaces, the desired singular values form one large cluster of zeros. Second, while convergence to the relevant invariant subspace is not necessary for low-rank approximation [6], such convergence is imperative to obtain a good null space approximation.
@@
Input: Matrix $A\in\mathbb{R}^{m\times n}$ with $m\geq n$ and $N$ zero singular values. Perturbation parameter $\epsilon>0$. Block size $d$. Number of iterations $\ell$.
@@ 2 Draw Gaussian random matrix $\Omega\in\mathbb{R}^{n\times d}$;
3 Perform block Lanczos to compute Lanczos decomposition $BZ_{\ell}=Z_{\ell}T_{\ell}+Q_{\ell+1}E_{\ell+1}$, where $Z_{\ell}$ is an orthonormal basis of Krylov subspace $\mathrm{span}\{\Omega,B\Omega,\dotsc,B^{\ell-1}\Omega\}$ and $T_{\ell}$ is block tridiagonal;
4 Compute an orthonormal basis $V_{Z}$ for the invariant space associated with $N$ smallest eigenvalues of $T_{\ell}$;
@@
Line 1 of Algorithm 1 extracts the null space approximation from Ritz vectors by computing the spectral decomposition of $T_{\ell}=Z_{\ell}^{\mathsf{T}}BZ_{\ell}$. Alternatively, one could remove the perturbation and compute the spectral decomposition of $Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}$ or, equivalently, the SVD of $AZ_{\ell}$. Preliminary numerical experiments indicate that this only yields minor accuracy improvements, at best, not justifying the extra computational effort needed to perform the alternative procedure.
The rest of this paper is organized as follows.
In Section 2, we analyze the impact of the random perturbation $\epsilon D$ introduced by Algorithm 1. In particular, Theorem 2.2 provides a new eigenvalue repulsion result for the perturbed zero eigenvalues of $A^{\mathsf{T}}A$. Our result differs from an existing result [19, Lemma 12] in two aspects: (1) By focusing on the repulsion of zero eigenvalues rather than the entire spectrum, we establish $\mathcal{O}(\epsilon)$ gaps instead of $\mathcal{O}(\epsilon^{2})$ gaps. (2) By considering relative eigenvalue gaps, we establish lower bound independent of the norm of $A$. At the same time, the perturbation incurs a limit on the attainable accuracy of the null space approximation, and we quantify this effect by a perturbation analysis.
Section 3 performs a self-contained convergence analysis of the randomized single-vector Lanczos method (corresponding to $d=1$ in Algorithm 1). While our analysis shares similarities with [19], we achieve more refined results by focusing on $d=1$ and considering the convergence of individual Ritz vectors. This analysis cumulates in Theorems 3.1 and 3.10, establishing bounds on the null space approximation, whose sharpness is verified by the numerical results.
In Section 4, we consider several practical improvements of Algorithm 1, including preconditioning, restarting, and partial reorthogonalization.
Assumptions and notation
Throughout this paper we assume that $A\in\mathbb{R}^{m\times n}$ satisfies $m\geq n$ and has $N\geq 1$ zero singular values. In accordance with Step 1 of Algorithm 1, we define
(2) $B:=A^{\mathsf{T}}A+\epsilon D,$
where $D$ is diagonal with nonnegative diagonal entries and $\epsilon>0$ controls the strength of the perturbation. We let $\lambda_{1},\ldots,\lambda_{n}$ denote the (nonnegative) eigenvalues of $B$, sorted such that
$0\leq\lambda_{1}\leq\dotsb\leq\lambda_{n}.$
[…]
Remark 1.2.
For null space computation, the cases $m\geq n$ and $m<n$ are fundamentally different because $m<n$ implies that $A$ has a null space of dimension at least $n-m$. When $n$ is much larger than $m$, say $n=2m$, one needs to resort to dense methods like the SVD, in the absence of additional structure.
@@ 2.1 Perturbation and repulsion of zero eigenvalues
We start by considering the impact of the perturbation on the $N$ zero eigenvalues of $A^{\mathsf{T}}A$. Suppose that $0<\epsilon\|D\|<\underline{\sigma}^{2}$, where $\underline{\sigma}$ denotes the smallest nonzero singular value of $A$. Then the zero and nonzero eigenvalues of $A^{\mathsf{T}}A$ will not be mixed by the perturbation $\epsilon D$. Indeed, using that $D$ is positive semi-definite, it follows from Weyl’s inequality [14, Theorem 4.3.1] that the eigenvalues $\lambda_{1},\dotsc,\lambda_{n}$ of $B$ satisfy
@@
Theorem 2.2 below shows that it is unlikely that the perturbed zero eigenvalues $\lambda_{1},\dotsc,\lambda_{N}$ remain very close to each other when $D$ is randomly chosen. To establish this eigenvalue repulsion result, we follow the proof of [19, Lemma 12] and make use of the following proposition.
@@
Consider the matrix $B$ defined in (2) and suppose that the diagonal entries of $D$ are drawn i.i.d. from the uniform distribution on $[0,1]$. Let $\underline{\sigma}$ denote the smallest nonzero singular value of $A$. If $0<\epsilon<\underline{\sigma}^{2}$, the inequality
@@
holds with probability at least $1-\delta$ for any $10<\delta<$ 1, where $C$ is a universal constant.
@@
Because of $\|D\|\leq 1$, it follows from (3) that $\lambda_{1},\dotsc,\lambda_{N}$ are contained in the interval $[0,\epsilon]$. We cover this interval with overlapping (smaller) intervals $\mathcal{I}_{j}=[j\gamma,(j+2)\gamma]$ for $j=0,\ldots,\epsilon/\gamma-2$, where—for the moment—$\gamma$ is an arbitrary number such that $20<\gamma\leq\epsilon/$ 2 and $\epsilon/\gamma$ is an integer. If two neighboring eigenvalues satisfy $\lambda_{i+1}-\lambda_{i}\leq\gamma$, then there clearly exists an interval $\mathcal{I}_{j}$ such that $\lambda_{i},\lambda_{i+1}\in\mathcal{I}_{j}$. Thus,
@@ 2.2 Perturbation of null space residual
The second result follows from a quadratic residual bound for eigenvalues in [17, Theorem 1]. To apply this bound, we let $V_{\perp}$ be such that $[V_{N},V_{\perp}]$ is an $n\times n$ orthogonal matrix and consider the gap
@@
Proposition 2.4 provides two kinds of perturbation bounds. The first bound does not require a gap but depends linearly on $\lVert B-B_{0}\rVert$, while the second one improves this to quadratic dependence at the expense of requiring a gap between $\lambda_{N}^{(0)}$ and $\lambda_{N+1}^{(0)}$.
@@
Let $V_{N}\in\mathbb{R}^{n\times N}$ be an orthonormal basis of the invariant subspace associated with the $N$ smallest eigenvalues of $B=A^{\mathsf{T}}A+\epsilon D$ for a symmetric positive semi-definite matrix $D$ with $\lVert D\rVert\leq 1$. If $20<\epsilon<\underline{\sigma}^{2}/$ 2, where $\underline{\sigma}$ denotes the smallest positive singular value of $A$, then
$\lVert AV_{N}\rVert\leq\epsilon(\underline{\sigma}^{2}-\epsilon)^{-1/2}\leq\sqrt{\epsilon}.$
@@
The result essentially follows from inserting $B_{0}=A^{\mathsf{T}}A$ into the second bound of Proposition 2.4.

### v2

# On a randomized small-block Lanczos method for large-scale null space computations

@@ Abstract
Computing the null space of a large sparse matrix $A$ is a challenging computational problem, especially if the nullity—the dimension of the null space—is not small. When applying a block Lanczos method to $A^{\mathsf{T}}A$ for this purpose, conventional wisdom suggests to use a block size $d$ that is not smaller than the nullity. In this work, we show how randomness can be utilized to allow for smaller $d$ without sacrificing convergence or reliability. Even $d=1$, corresponding to the standard single-vector Lanczos method, becomes a safe choice. This is achieved by using a small random diagonal perturbation, which moves the zero eigenvalues of $A^{\mathsf{T}}A$ away from each other, and a random initial guess. We analyze the effect of the perturbation on the attainable quality of the null space and derive convergence results that establish robust convergence for $d=1$. As demonstrated by our numerical experiments, a smaller block size combined with restarting and partial reorthogonalization results in reduced memory requirements and computational effort. It also allows for the incremental computation of the null space, without requiring a priori knowledge of the nullity. Our algorithm is best suited for situations when the nullity of $A$ is moderate.
@@ 1 Introduction
When $A$ is large and has no particular structure beyond being sparse, one needs to resort to Krylov subspace methods, which utilize matrix-vector products with $A$ and its transpose $A^{\mathsf{T}}$. Let $N$ denote the null space dimension of $A$. When $N=1$, standard choices include the Lanczos method for computing the smallest eigenvalue and eigenvector of $A^{\mathsf{T}}A$ or the mathematically equivalent, but numerically more reliable Golub–Kahan bidiagonalization directly applied to $A$ [11, Chapter 10]. When $N>1$, the Lanczos method (with a single starting vector) becomes unreliable and prone to miss components of the null space. In fact, in exact arithmetic only a one-dimensional subspace of the null space can be extracted via Lanczos [26, Section 12.2]. Block Lanczos methods address this issue by utilizing a block of $d>1$ starting vectors instead of a single starting vector [26, Section 13.10]. Common wisdom (see [26, Section 13.12] for examples) and existing convergence analyses all suggest to choose $d$ not smaller than the size of the eigenvalue cluster of interest. Applied to null space computation this would require $d\geq N$ and in exact arithmetic this condition is indeed necessary. The presence of round-off error blurs the situation. As an example, consider the $10064\times 10064$ diagonal matrix
(1) $A=\mathrm{diag}\Bigl(0,\dotsc,0,1+\frac{1}{10000},1+\frac{2}{10000},\dotsc,2\Bigr),$
that is, the first $64$ diagonal entries are zero. Because this matrix is symmetric positive semi-definite, one can apply block Lanczos directly to $A$ in order to compute its null space. When choosing a block of $d$ (Gaussian) random starting vectors, one would expect to obtain no more than $d$ (approximate) zero eigenvalues and corresponding eigenvectors. However, when executing block Lanczos with full reorthogonalization in double precision, limiting it to $1024$ matrix-vector multiplications with $A$ in total and declaring eigenvalues less than $10^{-4}$ as zero, we obtained the results reported in Table 1.
Table 1: Null space dimension obtained when applying block Lanczos with $d$ random starting vectors for 1024 matrix-vector multiplications to the matrix from Eq. 1.
Block size $d$ 1 2 4 8 16 32 64
Dimension 45 46 48 48 48 64 64
While the null space dimension estimated by block Lanczos with small block size $d$ is erratic and often smaller than $64$, it is always larger than $d$. In fact, already for $d=32$, the correct null space dimension is detected. Experiments with other matrices lead to similar findings, suggesting that round-off error breaks some of the curse incurred by eigenvalue clusters, but it does not fully address the convergence issues of single-vector or small-block Lanczos methods either.
@@
Inspired by [22], we propose a randomized small-block Lanczos method for null space calculation. Our method assumes that there is a significant gap between the zero and nonzero singular values. We first modify the matrix $A^{\mathsf{T}}A$ by adding a small random diagonal perturbation, which separates its zero eigenvalues. This separation enables us to use the small-block Lanczos method to capture an approximation for the whole null space basis of $A$. Compared to the task of low-rank approximation, there are two major differences. First, in low-rank approximation, the largest few singular values are usually unknown and, in most applications, they do not form tight large clusters. In the context of null spaces, the desired singular values form one large cluster of zeros. Second, while convergence to the relevant invariant subspace is not necessary for low-rank approximation [7], such convergence is imperative to obtain a good null space approximation.
The rest of this paper is organized as follows. In Section 2, we analyze the impact of the random perturbation on $A^{\mathsf{T}}A$. In particular, Theorem 2.2 provides a new eigenvalue repulsion result for the perturbed zero eigenvalues of $A^{\mathsf{T}}A$. Our result differs from an existing result [22, Lemma 12] in two aspects: (1) By focusing on the repulsion of zero eigenvalues rather than the entire spectrum, we establish $\mathcal{O}(\epsilon)$ gaps instead of $\mathcal{O}(\epsilon^{2})$ gaps. (2) By considering relative eigenvalue gaps, we establish a lower bound independent of the norm of $A$. At the same time, the perturbation incurs a limit on the attainable accuracy of the null space approximation, and we quantify this effect by a perturbation analysis.
Section 3 performs a self-contained convergence analysis of the randomized single-vector Lanczos method (applied to the perturbed matrix). While our analysis shares similarities with [22], we achieve more refined results by focusing on $d=1$ and considering the convergence of individual Ritz vectors. This analysis cumulates in Theorems 3.1 and 3.10, establishing bounds on the null space approximation, whose sharpness is verified by the numerical results.
In Section 4, we consider several practical improvements, including preconditioning, restarting, and partial reorthogonalization, culminating into Algorithm 2, our main algorithm. Note that, due to restarting, the convergence analysis of Section 3 does not apply to this algorithm. We also stress that our algorithm is conceptually quite close to existing restarted eigenvalue solvers, such as ARPACK [18], with the notable exception of the partial reorthogonalization strategy (which becomes important when dealing with larger null space dimensions).
1.1 Assumptions and notation
Throughout this paper we assume that $A\in\mathbb{R}^{m\times n}$ satisfies $m\geq n$ and has $N\geq 1$ zero singular values. We let $\underline{\sigma}$ denote the smallest positive singular value of $A$. We assume that there is a (significant) gap between $\underline{\sigma}$ and 0, which allows us to pick $\epsilon$ such that $0<\epsilon<\underline{\sigma}^{2}$. Such an assumption on the separation between zero and non-zero singular values is reasonable for matrices that have an exact null space in the absence of round-off error. Our method is not suitable when the singular values decay gradually to zero; see [3] for examples of such matrices.
Remark 1.1.
For null space computation, the cases $m\geq n$ and $m<n$ are fundamentally different because $m<n$ implies that $A$ has a null space of dimension at least nmn-m. Our algorithm can also be applied to this situation and our analysis carries over. However, when $n$ is much larger than $m$ (say, $n=2m$), its efficiency is compromised by the large null space dimension. In these situations, one needs to resort to dense methods like the SVD, unless there is additional structure that can be exploited.
Throughout the remainder of the paper, we use $\lVert\cdot\rVert$ to denote the Euclidean norm of a vector and the spectral norm of a matrix.
1.2 Basic algorithm
In accordance with the first step, we define
(2) $B:=A^{\mathsf{T}}A+\epsilon D,$
where $D$ is diagonal with nonnegative diagonal entries and $\epsilon>0$ controls the strength of the perturbation. The initial $\Omega\in\mathbb{R}^{n\times d}$ is chosen as a Gaussian random matrix, i.e., all entries are i.i.d. (independent and identically distributed) from the standard normal distribution $\mathcal{N}(0,1)$. In the second step, we perform block Lanczos with $\ell$ iterations to compute an orthonormal basis $Z_{\ell}$ for the Krylov subspace [^1] Note that our definition of a block Krylov subspace is different from some of the literature [14], but consistent with [36, 22].
$\mathcal{K}_{\ell}(B,\Omega):=\mathop{\mathrm{range}}[\Omega,B\Omega,\dotsc,B^{\ell-1}\Omega].$
To simplify the discussion, we will always assume that $\mathcal{K}_{\ell}(B,\Omega)$ has full rank and, therefore, $Z_{\ell}$ is of size $n\times\ell d$. For this subspace to contain any approximation to the $N$-dimensional null space, $\ell d$ must be at least $N$. Note that the block Lanczos method also returns the compression $T_{\ell}=Z_{\ell}^{\mathsf{T}}BZ_{\ell}\in\mathbb{R}^{\ell d\times\ell d}$. We use a Rayleigh–Ritz process to extract an approximate null space from $\mathcal{K}_{\ell}(B,\Omega)$. More specifically, the spectral decomposition of $T_{\ell}$ is computed to obtain the smallest $N$ eigenvalues of $T_{\ell}$ (called Ritz values) along with an orthonormal basis $V_{Z}$ of the corresponding eigenvectors. Then $V=Z_{\ell}V_{Z}$ contains the corresponding Ritz vectors, which tend to approximate eigenvectors belonging to the smallest eigenvalues of $B$. This makes it natural to choose $V$ as an approximate null space basis of $A$, leading to Algorithm 1.
@@ 1.2 Basic algorithm
Input: Matrix $A\in\mathbb{R}^{m\times n}$ with $m\geq n$ and NN zero singular values. Perturbation parameter $\epsilon>0$. Block size $d$. Number of iterations $\ell$ such that $\ell d\geq N$.
@@ 2 Draw Gaussian random matrix $\Omega\in\mathbb{R}^{n\times d}$;
3 Perform the block Lanczos method for $\ell$ iterations to compute an orthonormal basis $Z_{\ell}\in\mathbb{R}^{n\times\ell d}$ for $\mathcal{K}_{\ell}(B,\Omega)$ and the block tridiagonal matrix $T_{\ell}=Z_{\ell}^{\mathsf{T}}BZ_{\ell}\in\mathbb{R}^{\ell d\times\ell d}$;
4 Compute an orthonormal basis $V_{Z}\in\mathbb{R}^{\ell d\times N}$ of eigenvectors associated with NN smallest eigenvalues of $T_{\ell}$;
@@
Note that the matrix $B$ is never formed explicitly in Algorithm 1, but it is only accessed through matrix-vector products, which are carried out through matrix-vector products with the original matrix $A$ and $A^{\mathsf{T}}$: $Bx=A^{\mathsf{T}}(Ax)+\epsilon(Dx)$.
@@
As explained above, Line 4 of Algorithm 1 extracts a null space approximation via computing the spectral decomposition of $T_{\ell}=Z_{\ell}^{\mathsf{T}}BZ_{\ell}$. Alternatively, one could remove the perturbation and compute the spectral decomposition of $Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}$ or, equivalently, the SVD of $AZ_{\ell}$. Preliminary numerical experiments indicate that this only yields minor accuracy improvements, at best, not justifying the extra computational effort needed to perform the alternative procedure.
@@ 2.1 Perturbation and repulsion of zero eigenvalues
We start by considering the impact of the perturbation on the $N$ zero eigenvalues of $A^{\mathsf{T}}A$. Suppose that $0<\epsilon\|D\|<\underline{\sigma}^{2}$, where we recall that $\underline{\sigma}$ denotes the smallest nonzero singular value of $A$. Then the zero and nonzero eigenvalues of $A^{\mathsf{T}}A$ will not be mixed by the perturbation $\epsilon D$. Indeed, using that $D$ is positive semi-definite, it follows from Weyl’s inequality [16, Theorem 4.3.1] that the eigenvalues $\lambda_{1},\dotsc,\lambda_{n}$ of $B$ satisfy
@@
Theorem 2.2 below shows that it is unlikely that the perturbed zero eigenvalues $\lambda_{1},\dotsc,\lambda_{N}$ remain very close to each other when $D$ is randomly chosen. To establish this eigenvalue repulsion result, we follow the proof of [22, Lemma 12] and make use of the following proposition.
@@
Proposition 2.1 is a probabilistic result: For an arbitrary but fixed interval $\mathcal{I}$, it is very unlikely that more than one eigenvalue ends up in that interval, provided that the interval is sufficiently small, $\lvert\mathcal{I}\rvert\ll\epsilon/n$. In particular, for an eigenvalue $\lambda_{i}$ of $B_{0}+\epsilon D$ it is very unlikely that $\lambda_{i-1}$ or $\lambda_{i+1}$ are very close to $\lambda_{i}$. Theorem 2.2 below is a consequence of these considerations. In contrast, when $\lvert\mathcal{I}\rvert\gtrsim\epsilon/n$, the probability bound of Proposition 2.1 may be larger than $1$ and no conclusion can be drawn.
@@
Consider the matrix $B$ defined in (2) and suppose that the diagonal entries of $D$ are drawn i.i.d. from the uniform distribution on $[0,1]$. If $0<\epsilon<\underline{\sigma}^{2}$, the inequality
@@
holds with probability at least $1-\delta$ for any $0<\delta<1$, where $C$ is a universal constant.
@@
[…]
@@ 2.2 Perturbation of null space residual
The second result follows from a quadratic residual bound for eigenvalues in [20, Theorem 1]. To apply this bound, we let $V_{\perp}$ be such that $V_{N},V_{\perp}]$ is an $n\times n$ orthogonal matrix and consider the gap
@@
Lemma 2.4 provides two kinds of perturbation bounds. The first bound does not require a gap but depends linearly on $\lVert B-B_{0}\rVert$, while the second one improves this to quadratic dependence at the expense of requiring a gap between $\lambda_{N}^{(0)}$ and $\lambda_{N+1}^{(0)}$.
@@
Let $V_{N}\in\mathbb{R}^{n\times N}$ be an orthonormal basis of the invariant subspace associated with the $N$ smallest eigenvalues of $B=A^{\mathsf{T}}A+\epsilon D$ for a symmetric positive semi-definite matrix $D$ with $\lVert D\rVert\leq 1$. If $0<\epsilon<\underline{\sigma}^{2}$ then
$\lVert AV_{N}\rVert\leq\epsilon(\underline{\sigma}^{2}-\epsilon)^{-1/2}.$
Furthermore, if $0<\epsilon<\underline{\sigma}^{2}/2$, then $\lVert AV_{N}\rVert\leq\sqrt{\epsilon}$.
@@
The result essentially follows from inserting $B_{0}=A^{\mathsf{T}}A$ into the second bound of Lemma 2.4.